• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 15:32:41 on December23,2003

    Entry number 115048

    keyword=Beam shift with right Septum at 250 uA

    Right Septum at 250 uA, left is off. With the new installed steel beam pipe, the spot at the dump is shift by a little more than 1 cm, Almost one order of magnitude better than before.

    FIGURE 1